OpenStudy (anonymous):

How do I find the period for f(x)=cot(5x+5)+5 ? I thought it was 2pi/5, but apparently its wrong, can someone please explain ?

OpenStudy (anonymous):

\[a+b\cos(k(t-c)) \\period = \frac{2\pi}{k}\]

OpenStudy (anonymous):

factor the 5 out of 5x+5 then we have 5(x+1) period = 2pi/k k=5

OpenStudy (anonymous):

so, if you have this form \[a+b\cos(kt-kc)) \]then you need to factor the k out to get \[a+b\cos(k(t-c)) \] if you have this form \[a+b\cos(kt) \] this is just a special case of the other one, but c = 0 so in this situation we still hve period = 2pi/k

OpenStudy (anonymous):

understand @dnova21 ?

OpenStudy (anonymous):

Yes, thank you !

OpenStudy (anonymous):

you dont have to factor the k out. either way its the same thing for period.

OpenStudy (anonymous):

Oh okay :)

OpenStudy (unklerhaukus):

cot≠cos

OpenStudy (anonymous):

Ah I just caught that :S I found the answer to be pi/5 nonetheless, so when you are working with cot you do pi/B instead of 2pi/b right ?

OpenStudy (unklerhaukus):

yes , because tan(x), and cot(x) have a period of π,

OpenStudy (anonymous):

Oh now I get it thanks for clarifying :)

OpenStudy (anonymous):

sorry i read wrong